OpenStudy (anonymous):

prove induction n^2 < n! , for n>=4 , if it is easier n! > n^2 i didnt get far, the basis case is true and (k+1)!= (k+1)k! > ...

OpenStudy (watchmath):

I am sorry did you just delete your earlier question with my answer in it?

OpenStudy (anonymous):

oh im sorry

OpenStudy (anonymous):

this program is strange , i didnt see anything in it

OpenStudy (anonymous):

i apologize earlier, anything you have to say is helpful :)

OpenStudy (anonymous):

it doesnt have to be induction

OpenStudy (anonymous):

but i have a proof of my own, the basis case is true 4! > 4^2 (k+1)! = k!(k+1) > k^2 ( 2 + 1) = k^2( 1 + 1 + 1) = k^2 + k^2 + k^2 > k^2 + 2k + 1

OpenStudy (anonymous):

oh wait, that doesnt work

OpenStudy (anonymous):

not the most elegant proof, but it might work

OpenStudy (watchmath):

Ok, if you want we can use induction here we go. So we assume that k^2 < k! Notice that that k>1 is equivalent to 2k+1 < 3k It follows that k>= 4 we have 2k+1 < 3k < (k)! k Now using the assumption k^2 < k! we have k^2 +(2k+1) < k! +(2k+1) + k! +(k)! k=(k+1)! So (k+1)^2 < (k+1)!

OpenStudy (anonymous):

that looks, good,. one sec, what was your other proof . now im curious

OpenStudy (watchmath):

So you haven't read it?

OpenStudy (anonymous):

im reading it , one sec

OpenStudy (watchmath):

The statement \(n^2<n!\) is equivalent to \(n<(n-1)!\) Notice that for \(n\geq 4\) we have \((n-1)!\geq (n-2)(n-1)\) So it is enough to prove \((n-2)(n-1)>n\) The inequality is equivalent to \(n^2-3n+2>n\) \(n^2-4n+2>0\) \(n(n-4)+2>0\) which is clearly true for \(n\geq 4\).

OpenStudy (anonymous):

ok thank you

OpenStudy (watchmath):

Are you trying to use this for some statement about the series? if it is the case maybe it will be easier to prove the statement for some bigger n (not 4) to make it more obvious.

OpenStudy (anonymous):

thats clever, can you look over my proof the basis case is true 4! > 4^2 (k+1)! = k!(k+1) > k^2 ( 2 + 1) = k^2( 1 + 1 + 1) = k^2 + k^2 + k^2 > k^2 + 2k + 1

OpenStudy (anonymous):

not that i know of about serious

OpenStudy (anonymous):

because k>2 for k>=4

OpenStudy (anonymous):

oh but then i need to show k^2 > 2k ?

OpenStudy (watchmath):

correct, that is what I was going to say :)

OpenStudy (watchmath):

well actually you need 2k^2 > 2k+1

OpenStudy (watchmath):

Notice that \((k-1)^2 \geq 0\) So \(k^2-2k+1 \geq 0\) So \(k^2\geq 2k-1\) For \(k\geq 4, k^2 > 2\) So \(2k^2> k^2+2\geq (2k-1)+2 =2k +1\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can i ask a quick question. im helping a student who thinks that you can form a solid of revolution , bounded by y=x^2, y = x+2, and revolve it about the y axis

OpenStudy (watchmath):

Yes I think you can but there is some overlap between the solid form by the region on the right of the y-axis and the solid form by the region on the left side of the y-axis

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

keeps freezing on me